You run a coffee shop, where you currently sell 200 coffees per year at $4 each, and 100 pastries per year at $3 each. You also currently pay $200 per year to rent the building space, and you pay a barista $100 per year to serve the coffee and pastries.

You just paid a consultant $50 to assess ways in which you can increase revenue. Based on his advice, you are thinking of purchasing a sign spinner, currently priced at $125, to direct nearby foot traffic to your shop. You will also need to hire someone to operate the sign spinner, at an annual salary of $75 per year.

Based on market surveys, you expect that the sign spinner will enable you to sell 150 more coffees and 50 more pastries per year. The sign-spinning machine has a useful life of five years, and cannot be re-sold. Suppose you book the machine as an asset upon purchase, and declare an annual depreciation expense of $25 per year throughout its useful life. Your relevant marginal tax rate is 30%.

Assuming the relevant occ is 10%, what is the NPV of this investment (i.e., investing in the sign spinner)? There will be partial credit, where applicable. Round all interim work to at least 4 decimal places.
Answer #1

Incremental cash inflows is the additional sale of coffee and pastries.
Incremental cash outflows is the salary to be paid to person who would operate the sign spinner.
As consulting fees is sunk cost it is not incremental
As already rent and barista is being paid fixed cost, they would not come into picture

Hence,
Incremental Cash flows per year=(150*4 + 50*3-75-25)*(1-30%)+25=480
NPV = -125+480/1.1 + 480/1.1^2 + 480/1.1^3 + 480/1.1^4 + 480/1.1^5 = $1,694.58